3: The individual demand function of the consumer is MWTP = 12-1/2Q where Q is the quantity of a private good. How many units of the good does the consumer buy if the price is 6? Please show the result and the way you derived it.

4: What is the total willingness to pay for the number of units that correspond to a price of 6 (the number of units from question 3)? Show the results and the way you derived it.

5: For two consumers with MWTP = 12-1/2Q for a public good Q, what is the aggregate marginal willingness to pay at Q=0? Please show the result and the way you derived it.
Answer #1

3) MWTP = 12 - 0.5Q

If P = 6,

6 = 12 - 0.5Q

Q = 12

4) P = 12 - 0.5 * 12

P = 6

5) If there are 2 consumer, each willing to pay $6,

Aggregate marginal willingness to pay at Q = 0,

MWTP = 12 - 0.5 * 0 = 12 (addition of both consumer willingness to pay)